# Retrieval-Augmented Decision Transformer: External Memory for In-context RL

### v2

@@ D.1 Dark-Room
In addition, we also show the learning curves on Dark-Room $10\times 10$ over the entire training phase in Figure 15. We evaluate after every 25K updates and observe a steady improvement in the average performances with every evaluation.
(a) 80 Train Goals
(b) 20 Eval Goals
Figure 15: Average performances on Dark-Room 10$\times$10 over the course of training for (a) train and (b) test tasks. We train each agent for 100K steps and evaluate every 25K steps. Curves are averaged across the 80 train and 20 evaluation tasks, respectively. We report mean reward (+ 95% CI) over 3 seeds.
@@ D.1.1 Attention Map Analysis
What happens if an optimal trajectory is retrieved in context? In Figure 16, we showcase this example. The goal location is located at grid cell (4,6). The attention maps exhibit high attention scores for the state and the RTG at the end of the retrieved trajectory. We also observe high attention scores for the state similar to the current state and the action selected in that state. The agent initially imitates the actions in the context trajectory, but deviates further into the episode. Once the agent reaches the goal state, the attention scores for states and RTGs at the end of the trajectory reduce considerably, because the agent need not pay attention to the retrieved context any more.
Figure 16: Attention map analysis for an optimal context-trajectory on Dark-Room $10\times 10$. We plot the retrieved context trajectory (left), the corresponding attention map, and actual agent state (right), across timesteps (1, 5, 10). Queries (input trajectory) are on the y-axis and keys (context trajectory) on the x-axis. We highlight the sub-sequence in the context trajectory with the highest attention score (left). To improve readability, we mask-out attention scores below a certain threshold, and only provide labels for token that exhibit the highest attention scores. The agent imitates the context trajectory and successfully finds the goal.
What happens if a suboptimal trajectory is retrieved in Context? Similarly, we show the corresponding example in Figure 17. The goal location is again in grid cell (4,6). The retrieved context trajectory reaches the final state (9,5). Similar to Figure 16, the attention maps exhibit high attention scores for the last state and RTG for that state, as well as for a state at a similar timestep. Previously, RA-DT imitated the action, but in this situation the agent picks a different route, as the context trajectory does not lead to a successful outcome.
Figure 17: Attention map analysis for a suboptimal context-trajectory on Dark-Room $10\times 10$. The agent selects a different route than present in the suboptimal context trajectory and explores the environment.
This analysis suggests, that RA-DT can develop capabilities to either imitate a given positive experience or to behave differently than a given negative experience.
@@ D.1.2 Exploration Analysis
Delusions in RA-DT. Furthermore, we find that in some unsuccessful trials, the agent repeatedly performs the same suboptimal action sequences. Ortega et al. [2021] refer to such behaviour as delusions. In Figure 19, we illustrate two examples in which the agent suffers from delusions and does not recover until the end of the episode.
@@ D.2 Maze-Runner
While RA-DT outperforms competitors, we observe a considerable performance gap between train mazes and test mazes (0.65 vs. 0.4 reward, see Figure 20). This indicates that RA-DT struggles to solve difficult, unseen mazes. We believe that this gap is an artifact of the small pre-training distribution of 100 mazes, and be closed by increasing the number of pre-training mazes. Furthermore, increasing the number of ICL trials may also enhance the performance.
@@
Figure 20: Average performance on (a) 100 train and (b) 20 test mazes at end of training (100K steps). We evaluate each agent for 30 episodes and report mean reward (+ 95% CI) over 3 seeds.
@@
Figure 21: ICL on (a) 100 train and (b) 20 test mazes at end of training (100K steps). We evaluate each agent for 30 episodes and report mean reward (+ 95% CI) over 3 seeds.
@@ D.3 Meta-World
Generally, we observe that RA-DT outperforms competitors on the evaluation tasks in terms of average performance. However, on training task, the average performance of RA-DT is lower than of the vanilla DT. AD and DPT lack behind both methods. One potential reason is the RTG conditioning, which biases DT and RA-DT towards higher quality behaviour.
@@
Figure 22: Learning curves on (a) ML45 and (b) MT5 over the full training period (200K). We evaluate each agent for 30 episodes and report mean reward (+ 95% CI) over 3 seeds.
@@
Figure 23: ICL performance on (a) ML45 and (b) MT5 at end of training (200K steps). We evaluate each agent for 30 episodes and report mean reward (+ 95% CI) over 3 seeds.
In addition, we provide the average rewards and data-normalized scores in for the MT5 evaluation tasks in Table 3.
@@ D.4 DMControl
In Figures 24 and 25, we show the training curves across the entire training period (200K steps), and the corresponding ICL curves at the end of training for both DMC11 and DMC5.
@@
Figure 24: Average performance on (a) DMC11 and (b) DMC5 at end of training (200K steps). We evaluate each agent for 30 episodes and report mean reward (+ 95% CI) over 3 seeds.
@@
Figure 25: ICL performance on (a) DMC11 and (b) DMC5 at end of training (200K steps). We evaluate each agent for 30 episodes and report mean reward (+ 95% CI) over 3 seeds.
@@ D.5 Procgen
RA-DT constructs bursty sequences.. Building on work by Chan et al. [2022], Raparthy et al. [2023] identified trajectory burstiness as one important property for ICL to emerge on the Procgen benchmark. A given sequence is considered bursty, if it contains at least two trajectories from the same seed (or level). Consequently, the agent obtains relevant information that it can leverage to predict the next action. Therefore, we follow Raparthy et al. [2023] and always provide a trajectory from the same seed in the context of AD and DPT. Indeed, we observed that this improves performance, compared to not taking trajectory burstiness into account. Interestingly, we found that RA-DT retrieves trajectories from the same or similar seeds (seed accuracy of 8080~{}8080%), that is, RA-DT automatically constructs bursty sequences. This intuitively makes sense, as retrieval directly searches for the most relevant experiences (see Section 3.2.3). Therefore, for RA-DT, we do not provide additional information that indicates with which environment seed the trajectory was generated.
@@ E.1 Retrieval outperforms sampling of experiences
RA-DT is conditioned on sub-trajectories via cross-attention. By default, RA-DT leverages retrieval to search for relevant sub-trajectories for a given input sequence. Instead of retrieval, sub-trajectories can be sampled at random from the external memory. Therefore, we conduct an ablation in which we swap the retrieval mechanism with random sampling of sub-trajectories during training. This is to investigate the effect of relevance of retrieved sub-trajectories on learning performance. We apply random sampling only during training and use our regular retrieval during inference.
In Figure 6a, we show the ICL curves for training RA-DT with retrieved sub-trajectories, sub-trajectories sampled from the same task as the input sequence, and sub-trajectories sampled uniformly across all tasks. We find that training with retrieval outperforms both sampling variants. Uniform sampling results in poor ICL performance. A reason for this, is that context trajectories from a different goal location, are not relevant for predicting actions in the current sequences. As a result, the model ignores the given context during the training phase, and subsequently is unable to leverage it during inference. In contrast, sampling sub-trajectories from the same task as the input sequence results in better ICL performance, as the model learns to make use of the context trajectories. Nevertheless, using retrieval results in even better ICL performance, as sub-trajectories are not only relevant for the current task, but also similar to the current situation.
@@ E.2 Reweighting Mechanism
Figure 28: Effect of the Reweighting Mechanism. Average performances on Dark-Room 10$\times$10 over the course of training for (a) train and (b) test tasks.
In addition, we conduct a sensitivity analysis on the $\alpha$ parameter used in the re-weighting mechanism that determines how strongly the utility scores influences the final retrieval score. $\alpha=1$ is used both during training for task-based reweighing and during evaluation for return-based reweighting (see Section 3). In Figure 29, we vary $\alpha$ (a) during training, or (b) during evaluation, while keeping the other fixed. We find that RA-DT perform well for a range of values, but performance declines if no re-weighting is employed ($\alpha=0$).
@@
Figure 29: Sensitivity analysis on $\alpha$ parameter used in re-weighting mechanism of RA-DT on Dark-Room 10$\times$10.
@@ E.3 Retrieval Regularization
Providing the agent with too similar trajectories, can encourage it to adopt copying behaviour instead of generating high-reward actions. To mitigate this, we found it useful to regularize the retrieval using three strategies: deduplication, similarity cut-off, and query dropout. To evaluate their individual impact on ICL performance, we systematically removed each one from RA-DT in Figure 30.
We find that deduplication plays the most significant role in enhancing performance. One reason, why deduplication is effective, is because RL datasets contain many very similar trajectories. Removing overlapping trajectories altogether is therefore beneficial for learning. Notably, deduplication also reduces the index size, thereby speeding-up the search process. The effect of deduplication may vary depending on dataset characteristics, such as state-action coverage [Schweighofer et al., 2022].
@@
Figure 30: Effect of Retrieval Regularization. Average performances on Dark-Room 10$\times$10 over the course of training for (a) train and (b) test tasks.
@@ E.4 Query Construction & Sequence Aggregation
In RA-DT, we aggregate the hidden states of an input trajectory using mean aggregation of state tokens over the context length $C$ to obtain the $d_{r}$-dimensional query representation. It is, however, possible to use the hidden states of other tokens to construct the query. Therefore, we provide empirical evidence for this design choice in Figure 31a. We compare aggregating states, rewards, actions, returns-to-gos, all tokens, or only using the very last hidden state. Indeed, we find that aggregating state tokens gives the best results.
@@
Figure 31: Ablations on important components of RA-DT conducted on Dark-Room 10$\times$10. In (a) we investigate sequence aggregations to construct the query for retrieval. By default, we average state-tokens in the sequence (”mean, s”). In (b) we vary the placement of cross-attention layers in the DT. In (c) we vary the number of steps in-between retrievals during evaluation. We find that RA-DT delivers robust performance across settings.
@@ E.5 Placement of Cross-Attention Layers
Next, we investigate the effect of the placement of the cross-attention layers in RA-DT. In Figure 31b, we therefore vary the placement of cross-attention layers in RA-DT. By default, we use cross-attention after every self-attention layer. We find that other choice also provide good results. While placing the cross-attention at bottom layers tends to be beneficial, placing them only upper level layers tends to hurt performance.
@@ E.7 Effect of retrieval-augmentation on Training efficiency
In Figure 32, we compare domain-specific/agnostic RA-DT to the three considered baselines on Dark-Room across gridsizes $10\times 10$, $20\times 20$, and $40\times 20$. We find that the domain-specific variant of RA-DT attains minor training speed-ups on $10\times 10$ and trains almost 7$\times$ faster than baselines on the largest grid. The domain-agnostic variant of RA-DT, in contrast, exhibits slower training times on $10\times 10$, but also trains significantly faster on the largest grid. Note that the differences among the three grid-sizes in the number of samples processed per second of RA-DT stem from the difference in sequence lengths ($C=50$ for $10\times 10$, $C=100$ for $20\times 20$/$40\times 20$) and batch sizes ($B=128$ for $10\times 10$/$20\times 20$, $B=32$ for $40\times 20$).
The efficiency gains of RA-DT are a direct result of the shorter required sequence lengths. In contrast to the baselines, the computational requirements of RA-DT do not grow with the episode length of the environment. Additional speed-ups can be achieved for RA-DT by pre-computing the retrieved trajectories prior to training similar to Borgeaud et al. [2022]. We also want to highlight that all baselines use FlashAttention to speed-up the training times and to ensure a fair comparison. Consequently, the empirical evidence demonstrates that RA-DT does not only improve the down-stream performance in the environments, but is is also significantly faster to train (up to 7×7$\times$7 ×).
(a) Dark-Room 10$\times$10
(b) Dark-Room 20$\times$20
(c) Dark-Room 40$\times$20
Figure 32: Training efficiency for all considered methods on (a) Dark-Room 10$\times$10, (b) Dark-Room 20$\times$20, (c) Dark-Room 40$\times$20. We measure training efficiency in terms of the number of samples processed per second (higher is better). RA-DT achieves considerably speed-ups, in particular for larger grid-sizes.
@@ E.8 Effect of retrieval-augmentation on Inference efficiency
Retrieval-augmentation also incurs computational overhead during inference. Therefore, we study the effect of retrieval-augmentation on the infernce efficiency of RA-DT, similar to Appendix E.7. We measure inference efficiency in the number of environment interaction steps performed per second (higher is better). Note that this metric includes the environment latency. We average the inference efficiency metric across episodes to get a more robust estimate and discard the first episode to exclude compilation times. We conduct our analysis on an A100 GPU and use the same inference setup as described in Appendix C.
@@
In Figure 33, we report the inference efficiency for domain-specific/agnostic RA-DT and the considered baselines on Dark-Room. For RA-DT, we report the inference times with $t\in\{1,25\}$ where $t$ represents the number of interaction steps between retrievals. In Appendix E.6 we found that increasing $t$ only results in minor performance drops for RA-DT. For $t=1$ RA-DT exhibits slightly slower inference speeds compared to the baselines. In contrast, for $t=25$ there is no significant difference in inference speed between RA-DT and the baselines.
(a) Dark-Room 10$\times$10
(b) Dark-Room 20$\times$20
(c) Dark-Room 40$\times$20
Figure 33: Inference efficiency for all considered methods on (a) Dark-Room 10$\times$10, (b) Dark-Room 20$\times$20, (c) Dark-Room 40$\times$20. We measure inference efficiency in terms of the number of environment interaction steps performend per second (higher is better).
@@
Note that the inference speed is roughly the same across grid sizes. This suggests that the inference time is not yet dominated by the quadratic cost of self-attention for $B=1$ and the sequence lengths we consider in this analysis. To further support this, we run an ablation in which we compare the inference efficiency for all baselines with and without FlashAttention (see Figure 35) on Dark-Room $40\times 20$. Indeed, we observe a significant drop in inference speed when FlashAttention is disabled.
@@
Figure 34: Effect of FlashAttention on inference efficiency of AD, DPT and DT on Dark-Room $40\times 20$. Disabling FlashAttention results in a considerable drop in inference speed.
@@
To conclude this analysis, our findings indicate that while RA-DT is slightly slower when retrieving on every step, it achieves comparable inference speeds to the baselines when retrieving less frequently. Importantly, the retrieval mechanism in RA-DT enables access to the entirety of the experiences collected across all ICL trials. In contrast, the baselines can only access experiences from a limit set of the most recent episodes that are preserved in the context (2 in our experiments). If we were to provide more context episodes to the baselines, the quadratic complexity of self-attention would kick in (similar to Figure 32). The ability of RA-DT to access a much broader set of experiences may be a reason for its enhanced down-stream performance.
Figure 35: Effect of FlashAttention on inference efficiency of AD, DPT and DT on Dark-Room $40\times 20$. Disabling FlashAttention results in a considerable drop in inference speed.

### v3

@@ D.1 Dark-Room
In addition, we also show the learning curves on Dark-Room $10\times 10$ over the entire training phase in Figure 15. We evaluate after every 25K updates and observe a steady improvement in the average performance with every evaluation.
@@ D.1.1 Attention Map Analysis
What happens if an optimal trajectory is retrieved in context? In Figure 16, we showcase this example. The goal location is located at grid cell (4,6). The attention maps exhibit high attention scores for the state and the RTG at the end of the retrieved trajectory. We also observe high attention scores for the state, similar to the current state and the action selected in that state. The agent initially imitates the actions in the context trajectory, but deviates further into the episode. Once the agent reaches the goal state, the attention scores for states and RTGs at the end of the trajectory reduce considerably, because the agent needs not pay attention to the retrieved context any more.
What happens if a suboptimal trajectory is retrieved in Context? Similarly, we show the corresponding example in Figure 17. The goal location is again in grid cell (4,6). The retrieved context trajectory reaches the final state (9,5). Similar to Figure 16, the attention maps exhibit high attention scores for the last state and RTG for that state, as well as for a state at a similar timestep. Previously, RA-DT imitated the action, but in this situation, the agent picks a different route, as the context trajectory does not lead to a successful outcome.
This analysis suggests that RA-DT can develop capabilities to either imitate a given positive experience or to behave differently than a given negative experience.
@@ D.1.2 Exploration Analysis
Delusions in RA-DT. Furthermore, we find that in some unsuccessful trials, the agent repeatedly performs the same suboptimal action sequences. Ortega et al. (2021) refer to such behaviour as delusions. In Figure 19, we illustrate two examples in which the agent suffers from delusions and does not recover until the end of the episode.
@@ D.2 Maze-Runner
While RA-DT outperforms competitors, we observe a considerable performance gap between train mazes and test mazes (0.65 vs. 0.4 reward, see Figure 20). This indicates that RA-DT struggles to solve difficult, unseen mazes. We believe that this gap is an artifact of the small pre-training distribution of 100 mazes and can be closed by increasing the number of pre-training mazes. Furthermore, increasing the number of ICL trials may also enhance the performance.
@@
Figure 20: Average performance on (a) 100 train and (b) 20 test mazes at end of training (100K steps). We evaluate each agent for 30 episodes and report the mean reward (+ 95% CI) over 3 seeds.
@@
Figure 21: ICL on (a) 100 train and (b) 20 test mazes at end of training (100K steps). We evaluate each agent for 30 episodes and report the mean reward (+ 95% CI) over 3 seeds.
@@ D.3 Meta-World
Generally, we observe that RA-DT outperforms competitors on the evaluation tasks in terms of average performance. However, on the training task, the average performance of RA-DT is lower than that of vanilla DT. AD and DPT lag behind both methods. One potential reason is the RTG conditioning, which biases DT and RA-DT towards higher quality behaviour.
@@
Figure 22: Learning curves on (a) ML45 and (b) MT5 over the full training period (200K). We evaluate each agent for 30 episodes and report the mean reward (+ 95% CI) over 3 seeds.
@@
Figure 23: ICL performance on (a) ML45 and (b) MT5 at end of training (200K steps). We evaluate each agent for 30 episodes and report the mean reward (+ 95% CI) over 3 seeds.
In addition, we provide the average rewards and data-normalized scores for the MT5 evaluation tasks in Table 3.
@@ D.4 DMControl
In Figures 24 and 25, we show the training curves across the entire training period (200K steps) and the corresponding ICL curves at the end of training for both DMC11 and DMC5.
@@
Figure 24: Average performance on (a) DMC11 and (b) DMC5 at end of training (200K steps). We evaluate each agent for 30 episodes and report the mean reward (+ 95% CI) over 3 seeds.
@@
Figure 25: ICL performance on (a) DMC11 and (b) DMC5 at end of training (200K steps). We evaluate each agent for 30 episodes and report the mean reward (+ 95% CI) over 3 seeds.
@@ D.5 Procgen
RA-DT constructs bursty sequences.. Building on work by Chan et al. (2022), Raparthy et al. (2023) identified trajectory burstiness as one important property for ICL to emerge on the Procgen benchmark. A given sequence is considered bursty if it contains at least two trajectories from the same seed (or level). Consequently, the agent obtains relevant information that it can leverage to predict the next action. Therefore, we follow Raparthy et al. (2023) and always provide a trajectory from the same seed in the context of AD and DPT. Indeed, we observed that this improves performance, compared to not taking trajectory burstiness into account. Interestingly, we found that RA-DT retrieves trajectories from the same or similar seeds (seed accuracy of 80~80% This intuitively makes sense, as retrieval directly searches for the most relevant experiences (see Section 3.2.3). Therefore, for RA-DT, we do not provide additional information that indicates with which environment the trajectory was generated.
@@ E.1 Retrieval outperforms sampling of experiences
RA-DT is conditioned on sub-trajectories via cross-attention. By default, RA-DT leverages retrieval to search for relevant sub-trajectories for a given input sequence. Instead of retrieval, sub-trajectories can be sampled at random from the external memory. Therefore, we conduct an ablation in which we swap the retrieval mechanism with random sampling of sub-trajectories during training. This is to investigate the effect of the relevance of retrieved sub-trajectories on learning performance. We apply random sampling only during training and use our regular retrieval during inference.
In Figure 6a, we show the ICL curves for training RA-DT with retrieved sub-trajectories, sub-trajectories sampled from the same task as the input sequence, and sub-trajectories sampled uniformly across all tasks. We find that training with retrieval outperforms both sampling variants. Uniform sampling results in poor ICL performance. A reason for this is that context trajectories from a different goal location are not relevant for predicting actions in the current sequences. As a result, the model ignores the given context during the training phase and subsequently is unable to leverage it during inference. In contrast, sampling sub-trajectories from the same task as the input sequence results in better ICL performance, as the model learns to make use of the context trajectories. Nevertheless, using retrieval results in even better ICL performance, as sub-trajectories are not only relevant for the current task, but also similar to the current situation.
@@ E.2 Reweighting Mechanism
Figure 28: Effect of the Reweighting Mechanism. Average performances on Dark-Room $\times10$ over the course of training for (a) train and (b) test tasks.
In addition, we conduct a sensitivity analysis on the $\alpha$ parameter used in the re-weighting mechanism that determines how strongly the utility scores influence the final retrieval score. $\alpha=1$ is used both during training for task-based reweighing and during evaluation for return-based reweighting (see Section 3). In Figure 29, we vary $\alpha$ (a) during training, or (b) during evaluation, while keeping the other fixed. We find that RA-DT performs well for a range of values, but performance declines if no re-weighting is employed ($\alpha=0$).
@@
Figure 29: Sensitivity analysis on $\alpha$ parameter used in re-weighting mechanism of RA-DT on Dark-Room $\times10$.
@@ E.3 Retrieval Regularization
Providing the agent with too similar trajectories can encourage it to adopt copying behaviour instead of generating high-reward actions. To mitigate this, we found it useful to regularize the retrieval using three strategies: deduplication, similarity cut-off, and query dropout. To evaluate their impact on ICL performance, we systematically removed each one from RA-DT in Figure 30.
We find that deduplication plays the most significant role in enhancing performance. One reason why deduplication is effective is that RL datasets contain many very similar trajectories. Removing overlapping trajectories altogether is therefore beneficial for learning. Notably, deduplication also reduces the index size, thereby speeding up the search process. The effect of deduplication may vary depending on dataset characteristics, such as state-action coverage (Schweighofer et al., 2022).
@@
Figure 30: Effect of Retrieval Regularization. Average performances on Dark-Room $\times10$ over the course of training for (a) train and (b) test tasks.
@@ E.4 Query Construction & Sequence Aggregation
In RA-DT, we aggregate the hidden states of an input trajectory using mean aggregation of state tokens over the context length $C$ to obtain the $d_{r}$-dimensional query representation. It is, however, possible to use the hidden states of other tokens to construct the query. Therefore, we provide empirical evidence for this design choice in Figure 31a. We compare aggregating states, rewards, actions, return-to-gos, all tokens, or only using the very last hidden state. Indeed, we find that aggregating state tokens gives the best results.
@@
Figure 31: Ablations on important components of RA-DT conducted on Dark-Room $\times10$. In (a) we investigate sequence aggregations to construct the query for retrieval. By default, we average state-tokens in the sequence (”mean, s”). In (b) we vary the placement of cross-attention layers in the DT. In (c) we vary the number of steps in-between retrievals during evaluation. We find that RA-DT delivers robust performance across settings.
@@ E.5 Placement of Cross-Attention Layers
Next, we investigate the effect of the placement of the cross-attention layers in RA-DT. By default, we use cross-attention after every self-attention layer. As a result, RA-DT has slightly more parameters than AD/DPT/DT (14M vs. 16.5M). Note that most of the parameters reside in the MLPs rather than in the attention layers. To verify that the increased parameter count is not responsible for the improved performance, we conduct an ablation study in Figure 31b, in which we vary the placement of the cross-attention layers. Interestingly, using a single cross-attention block at the first layer yields similar performance to using cross-attention after every layer. Moreover, while placing the cross-attention layers at the bottom layers tends to be beneficial, placing them only upper-level layers tends to hurt performance.
@@ E.7 Effect of retrieval-augmentation on Training efficiency
[…]
The efficiency gains of RA-DT are a direct result of the shorter required sequence lengths. In contrast to the baselines, the computational requirements of RA-DT do not grow with the episode length of the environment. Additional speed-ups can be achieved for RA-DT by pre-computing the retrieved trajectories prior to training, similar to Borgeaud et al. (2022). We also want to highlight that all baselines use FlashAttention to speed up the training times and to ensure a fair comparison. Consequently, the empirical evidence demonstrates that RA-DT not only improves the downstream performance in the environments, but is also significantly faster to train (up to $7\times$).
(a) Dark-Room $\times10$
(b) Dark-Room $\times20$
[…]
@@ E.8 Effect of retrieval-augmentation on Inference efficiency
Retrieval-augmentation also incurs computational overhead during inference. Therefore, we study the effect of retrieval-augmentation on the inference efficiency of RA-DT, similar to Appendix E.7. We measure inference efficiency in the number of environment interaction steps performed per second (higher is better). Note that this metric includes the environment latency. We average the inference efficiency metric across episodes to get a more robust estimate and discard the first episode to exclude compilation times. We conduct our analysis on an A100 GPU and use the same inference setup as described in Appendix C.
@@
In Figure 33, we report the inference efficiency for domain-specific/agnostic RA-DT and the considered baselines on Dark-Room. For RA-DT, we report the inference times with $t\in\{1,25\}$ where $t$ represents the number of interaction steps between retrievals. In Appendix E.6, we found that increasing $t$ only results in minor performance drops for RA-DT. For $t=1$ RA-DT exhibits slightly slower inference speeds compared to the baselines. In contrast, for $t=25$ there is no significant difference in inference speed between RA-DT and the baselines.
(a) Dark-Room $\times10$
(b) Dark-Room $\times20$
[…]
@@
Note that the inference speed is roughly the same across grid sizes and consequently episode lengths. This suggests that the inference time is not yet dominated by the quadratic cost of self-attention for $B=1$ and the sequence lengths we consider in this analysis. This is because inference remains memory-bound with $B=1$ when using FlashAttention on the hardware setup we consider. For environments with even longer episode lengths, higher batch sizes, or bigger models, the inference would move towards a compute-bound regime, and similar speed-ups for RA-DT are to be expected as for training time. To further support this, we run an ablation in which we compare the inference efficiency for all baselines with and without FlashAttention (see Figure 34) on Dark-Room $40\times 20$. Indeed, we observe a significant drop in inference speed when FlashAttention is disabled.
@@
Figure 34: Effect of FlashAttention on inference efficiency of AD, DPT, and DT on Dark-Room $40\times 20$. Disabling FlashAttention results in a considerable drop in inference speed.
@@
To conclude this analysis, our findings indicate that while RA-DT is slightly slower when retrieving on every step, it achieves comparable inference speeds to the baselines when retrieving less frequently. Importantly, the retrieval mechanism in RA-DT enables access to the entirety of the experiences collected across all ICL trials. In contrast, the baselines can only access experiences from a limited set of the most recent episodes that are preserved in the context (2 in our experiments). If we were to provide more context episodes to the baselines, the quadratic complexity of self-attention would kick in (similar to Figure 32). The ability of RA-DT to access a much broader set of experiences may be a reason for its enhanced downstream performance.